Calculate the free energy change for the reaction shown below at 25oC. 2NO + O2 → NO3

1
Expert's answer
2022-05-23T17:01:05-0400

"\Delta G=\Delta H-T\Delta S"

"2NO+O_2\rightarrow 2NO_2"

[2ΔHf(NO2 (g))] - [2ΔHf(NO (g)) + 1ΔHf(O2 (g))]

[2(33.18)] - [2(90.25) + 1(0)] = -114.14 kJ

[2ΔSf(NO2(g))] - [2ΔSf(NO (g)) + 1ΔSf(O2 (g))]

[2(239.95)] - [2(210.65) + 1(205.03)] = -146.43 J/K

T=25+273=298K

[2ΔGf(NO2 (g))] - [2ΔGf(NO (g)) + 1ΔGf(O2 (g))]

[2(51.3)] - [2(86.57) + 1(0)] = -70.54 kJ

-70.54 kJ (spontaneous)